given: the line containing the points (1,-6) and (5,3) is perpendicular to the containing the points (4,8) and (-3,K).
1. find the value of K.
2. find the equation of each line in the form of y=mx+b.
could someone please help me with this,I am messing something up. Thank you for your help.

Answer by stanbon(48558)

given: the line containing the points (1,-6) and (5,3) is perpendicular to the containing the points (4,8) and (-3,K).
1. find the value of K.
2. find the equation of each line in the form of y=mx+b
#1 Equation of the 1st line:
m=[3+6]/[5-1]= 9/4
Using the point (5,3) and the form y=mx+b, substitute to get:
3=(9/4)(5)+b
b=12/4 - 45/4 = -33/4
Equation of that line is y=(9/4)x-33/4
#2 Since perpendicular, this line must have slope = -4/9
From points you get m=[K-8]/[-7]=-4/9
Then K=28/9+8
K=28/9+72/9= 100/9
Now get equation of 2nd line:
Using point(4,8) get 8=(-4/9)(4)+b
b=72/9+16/9 = 88/9
Equation of the line is y=-4/9x+88/9
Cheers,
stan H.